# Starling: An I/O-Efficient Disk-Resident Graph Index Framework for High-Dimensional Vector Similarity Search on Data Segment

### 04 Jan 2024 v1

@@ Abstract.
High-dimensional vector similarity search (HVSS) is receiving a spotlight as a powerful tool for various data science and AI applications. As vector data grows larger, in-memory indexes become extremely expensive because they necessitate substantial expansion of main memory resources. One possible solution is to use disk-based implementation, which stores and searches vector data in high-performance devices like NVMe SSDs. However, HVSS for data segments is still challenging in vector databases, where one machine has multiple segments for system features (like scaling) purposes. In this setting, each segment has limited memory and disk space, so HVSS on the data segment needs to balance accuracy, efficiency, and space cost. Existing disk-based methods are sub-optimal because they do not consider all these requirements together.
In this paper, we present $\mathtt{Starling}$, an I/O-efficient disk-resident graph index framework that optimizes data layout and search strategy in the segment. It has two main components: (1) a data layout that includes an in-memory navigation graph and a reordered disk-based graph with locality enhancement, which reduces the search path length and disk bandwidth wastage; and (2) a block search strategy that minimizes expensive disk I/Os when executing a vector query. We conduct extensive experiments to verify $\mathtt{Starling}$’s effectiveness, efficiency, and scalability. On a data segment with 2GB memory and 10GB disk capacity, $\mathtt{Starling}$ can maintain up to 33 million vectors in 128 dimensions, and serve HVSS with more than 0.9 average precision and top-10 recall rate, and latency of under 1 millisecond. The results show that $\mathtt{Starling}$ exhibits 43.9$\times$ higher throughput with 98% lower query latency than state-of-the-art methods under the same accuracy.
@@ Acknowledgements.
($\Rightarrow$): If there is a block shuffling of $G$ with $OR(G)=1$, then each block contains three cliques, and the sum of their sizes is $\Omega$. We can map each clique size in a block to an integer and get a solution to the triple shuffling problem.
($\Leftarrow$): If there is a solution to the triple shuffling problem, then the integers can be grouped into triples that sum up to $\Omega$. We can map each triple to a block and each integer to a clique size in that block, resulting in a block shuffling of $G$ with $OR(G)=1$.
@@
Suppose that we have a polynomial time approximation algorithm with a finite approximation factor for the block shuffling problem. We can use it to solve the triple shuffling problem. Case 1: If there is a block shuffling of $G$ with $OR(G)=1$, then the triple shuffling problem has a solution. Case 2: If there is no block shuffling of $G$ with $OR(G)=1$, then $OR(G)$ is less than 1 and the triple shuffling problem has no solution. Any approximation algorithm with a finite approximation factor for the block shuffling problem must be able to tell these two cases apart to solve the triple shuffling problem. Since we know that checking whether it returns $OR(G)=1$ or not is NP-hard, which contradicts the assumption.

### 02 Mar 2024 v3

@@ Abstract.
High-dimensional vector similarity search (HVSS) is gaining prominence as a powerful tool for various data science and AI applications. As vector data scales up, in-memory indexes pose a significant challenge due to the substantial increase in main memory requirements. A potential solution involves leveraging disk-based implementation, which stores and searches vector data on high-performance devices like NVMe SSDs. However, implementing HVSS for data segments proves to be intricate in vector databases where a single machine comprises multiple segments for system scalability. In this context, each segment operates with limited memory and disk space, necessitating a delicate balance between accuracy, efficiency, and space cost. Existing disk-based methods fall short as they do not holistically address all these requirements simultaneously.
In this paper, we present $\mathtt{Starling}$, an I/O-efficient disk-resident graph index framework that optimizes data layout and search strategy within the segment. It has two primary components: (1) a data layout incorporating an in-memory navigation graph and a reordered disk-based graph with enhanced locality, reducing the search path length and minimizing disk bandwidth wastage; and (2) a block search strategy designed to minimize costly disk I/O operations during vector query execution. Through extensive experiments, we validate the effectiveness, efficiency, and scalability of $\mathtt{Starling}$. On a data segment with 2GB memory and 10GB disk capacity, $\mathtt{Starling}$ can accommodate up to 33 million vectors in 128 dimensions, offering HVSS with over 0.9 average precision and top-10 recall rate, and latency under 1 millisecond. The results showcase $\mathtt{Starling}$’s superior performance, exhibiting 43.9$\times$ higher throughput with 98% lower query latency compared to state-of-the-art methods while maintaining the same level of accuracy.
@@ Acknowledgements.
($\Rightarrow$): If there is a block shuffling of $G$ with the largest $OR(G)$ (let $c$ be the largest $OR(G)$), then each block contains three cliques, and the sum of their sizes is $\Omega$. We can map each clique size in a block to an integer and get a solution to the triple shuffling problem.
($\Leftarrow$): If there is a solution to the triple shuffling problem, then the integers can be grouped into triples that sum up to $\Omega$. We can map each triple to a block and each integer to a clique size in that block, resulting in a block shuffling of $G$ with $OR(G)=c$.
@@
Suppose that we have a polynomial time approximation algorithm with a finite approximation factor for the block shuffling problem. We can use it to solve the triple shuffling problem. Case 1: If there is a block shuffling of $G$ with the largest $OR(G)$, then the triple shuffling problem has a solution. Case 2: If there is no block shuffling of $G$ with the largest $OR(G)$, then $OR(G)$ is less than $c$ and the triple shuffling problem has no solution. Any approximation algorithm with a finite approximation factor for the block shuffling problem must be able to tell these two cases apart to solve the triple shuffling problem. Since we know that checking whether it returns the largest $OR(G)$ or not is NP-hard, which contradicts the assumption.